OpenStudy (anonymous):

what is the solution set for the following equations ? x = -y and x+2y=6? is it : A. -2,2 B. 2, -2 iC. 6,-6 D. -6,6

OpenStudy (anonymous):

what is this for?

OpenStudy (anonymous):

algebra

OpenStudy (anonymous):

thought it would be C but not sure? do you know how to do this?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

do you know how to do substitution?

OpenStudy (anonymous):

kind of but not really...just tell me which choice it would be and then i will worry about the work . i just need to get this answer right . this hw assignment is for a grade and i got a bad grade in math and need to bring it up

OpenStudy (anonymous):

how about i tell you how to get there, and you worry about the answer

OpenStudy (anonymous):

i tried substitution just now but it didnt get me anywhere

OpenStudy (anonymous):

just explain the answer and how you got there

OpenStudy (anonymous):

So substitute x = -y into your other equation. Then you get the value for y. Substitute the value of y in either equation.

OpenStudy (anonymous):

i did !

OpenStudy (anonymous):

i got x=1/2x-3 and y=-1/2x+3

OpenStudy (anonymous):

If x=-y you plug =y in for x so- -y+2y=6 y=6 so x=-(6) the right answer is C, but im not sure why C and D are the same thing just opposite?

OpenStudy (anonymous):

thanks @pinkberry501 !

OpenStudy (anonymous):

that's how you would do it^

OpenStudy (anonymous):

also had 1 more question : the graph of which equation does NOT pass through the origin? A. y=x B. y= -x C. y=0 D. y=1

OpenStudy (anonymous):

since im on the same topic

OpenStudy (anonymous):

y=1 doesn't pass through the origin

OpenStudy (anonymous):

thanks so much! you are the best!!

OpenStudy (anonymous):

thats what i thought

OpenStudy (anonymous):

haha no problem!

OpenStudy (anonymous):

can i ask u one more question?

OpenStudy (anonymous):

Sure!

OpenStudy (anonymous):

which equation of the line that is parallel to y=2x-8 passes through the point (0, -3) A. y= 2x+3 B. y=2x-3 C. y=-1/2x+3 D. y=-1/2x-3

OpenStudy (anonymous):

i said D but im not sure

OpenStudy (anonymous):

what grade r u in anway by the way?

OpenStudy (anonymous):

well you would use y=mx+b -3=2(0)+b and if the line is parallel the slope would be 2 so i would say B. -3=0+b You did the opposite reciprocal of the slope (2) which you do -3=b for a perpendicular line.

OpenStudy (anonymous):

oops that got messed up lol & im in 9th grade

OpenStudy (anonymous):

same!

OpenStudy (anonymous):

but so what answer would it be ?

OpenStudy (anonymous):

The answer would be B and oh nice hahaa are you in algebra 2?

OpenStudy (anonymous):

nope Algebra 1B

OpenStudy (anonymous):

i will tag u in quetions i have since ur smart !

OpenStudy (anonymous):

ur a great helper

OpenStudy (anonymous):

Haha thanks! I'm in geometry so I could probably help you with most of the questions

OpenStudy (anonymous):

when the expressions x^2 -9 and x^2 - 5x+6 are factored, a common factor is: A. x+3 B. x-3 (i think but not sure) C. x-2 D. x^2

OpenStudy (anonymous):

@pinkberry501

OpenStudy (anonymous):

well x^2-9 =(x+3)(x-3) and x^2-5x+6= (x-2)(x-3) so yup you're right.

OpenStudy (anonymous):

wow! thanks again!